OpenStudy (anonymous):

solve elimination to solve linear system 3x+2y=-5 4x-3y=16

OpenStudy (zarkon):

multiply the first equation by 3 and the 2nd equation by 2 and add them

OpenStudy (zarkon):

9x+6y=-15 8x-6y=32 ---------- 17x+0y=17 x=1